# Bias Similarity Across Large Language Models

### v3

@@ Abstract
Bias in Large Language Models remains a critical concern as these systems are increasingly deployed in high-stakes applications. Yet most fairness evaluations rely on scalar metrics or single-model analysis, overlooking how biases align—or diverge—across model families, scales, and tuning strategies. In this work, we reframe bias similarity as a form of functional similarity and evaluate 24 LLMs from four major families on over one million structured prompts spanning four bias dimensions. Our findings uncover that fairness is not strongly determined by model size, architecture, instruction tuning, or openness. Instead, bias behaviors are highly context-dependent and structurally persistent, often resistant to current alignment techniques. Contrary to common assumptions, we find that open-source models frequently match or outperform proprietary models in both fairness and utility. These results call into question the default reliance on proprietary systems and highlight the need for behaviorally grounded, model-specific audits to better understand how bias manifests and endures across the LLM landscape.
Bias Similarity Across Large Language Models
Hyejun Jeong UMass Amherst hjeong@umass.edu                        Shiqing Ma UMass Amherst shiqingma@umass.edu                        Amir Houmansadr UMass Amherst amir@cs.umass.edu
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used in sensitive domains such as education, hiring, healthcare, and law. However, these systems often exhibit undesirable social biases—amplifying stereotypes, reinforcing inequality, or generating unsafe content Ferrara (2023); Sweeney (2013). A growing body of work has identified such behaviors, but understanding how these biases vary across models remains underexplored.
Prior studies often assess bias within individual models, typically using scalar metrics like accuracy Oketunji et al. (2023); Parrish et al. (2021) or bias scores Nadeem et al. (2020). Despite its convenience, these often obscure how models behave across demographic groups. While recent works have begun examining functional similarity—e.g., comparing model outputs or decision patterns Klabunde et al. (2023b); Li et al. (2021); Guan et al. (2022)—these efforts rarely center on fairness behavior or evaluate models beyond the open-source landscape.
In this work, we reframe bias similarity as a form of functional similarity: do different LLMs exhibit similar bias behaviors across the same prompts? This reframing shifts the focus from “Is a model fair?” to Which models behave similarly with respect to fairness?—a perspective that helps uncover underlying behavioral structures and identifies which factors (e.g., size, tuning, architecture) drive bias alignment.
To that end, we evaluate 24 LLMs from four major families (LLaMA, Gemma, GPT, and Gemini), ranging from 7B to 70B parameters. We analyze over one million structured prompts from BBQ Parrish et al. (2021) and UnQover Li et al. (2020), as well as open-ended completions from StereoSet Nadeem et al. (2020).
Our key findings reveal:
  • Fairness is multi-dimensional and model-specific: Performance on fairness varies significantly across bias dimensions and models—even within the same family—revealing that fairness cannot be meaningfully captured by a single metric or model-wide generalization.
  • Proprietary models prioritize caution at the cost of utility: These models frequently default to “unknown,” reducing potentially biased outputs but often failing to provide informative or correct answer to sensitive prompts—even when the context allows for a specific answer.
  • Instruction tuning may mask bias through abstention: While tuning encourages conservative behavior (e.g., abstaining on ambiguous prompts), it often fails to eliminate directional bias in forced-choice settings, reflecting a shallow form of fairness.
  • Open-source models can be competitive: Open-source LLMs demonstrate fairness performance comparable to, and in some cases exceeding, proprietary models. Their strong performance with transparency in training and architecture, challenges the assumption that larger, closed models are inherently fairer.
These findings underscore the importance of comprehensive, context-aware fairness evaluations and the need for ongoing research into robust mitigation strategies that address structural bias in LLMs.
@@ 2 Related Works
We summarize prior work in two areas: bias assessment in LLMs and methods for identifying similarity across models.
@@ 2.1 Bias Assessment in LLMs
Numerous studies show that LLMs exhibit social biases across dimensions such as gender, race, religion, and socioeconomic status. In response, datasets like StereoSet Nadeem et al. (2020), CrowS-Pairs Nangia et al. (2020), UnQover Li et al. (2020), and BBQ Parrish et al. (2021) have been developed to quantify these biases through masked completions or multiple-choice QA.
Bias is variably defined: as systemic disparities across groups Manvi et al. (2024), skewed performance across demographics Oketunji et al. (2023); Gupta et al. (2023), representational harms via stereotyping Lin et al. (2024); Zhao et al. (2023), or outcomes rooted in power imbalances Gallegos et al. (2024). Yet, distinguishing biased from factual generalizations remains challenging. For example, answering “younger people” to a question about tech adaptability may be statistically grounded but still propagate age-related stereotypes Vaportzis et al. (2017).
Recent studies have used LLMs themselves as evaluators Ye et al. (2024); Shi et al. (2024), though this raise concerns about evaluator inconsistency and model-induced bias eading to unreliable assessments Stureborg et al. (2024). Others analyzed how bias manifests across architectures Yeh et al. (2023) or in generation tasks via pairwise comparison Liusie et al. (2023), stereotype frequency Bahrami et al. (2024), and retrieval-based exposure Dai et al. (2024). Still, most focused on individual models and did not assess whether bias patterns generalize across architectures, sizes, or tuning strategies.
Our work shifts focus to cross-model behavioral comparison, investigating whether bias patterns persist or diverge across family, open- and closed-source models, and tuning strategies.
@@ 2.2 LLM Similarity and Behavioral Alignment
Prior work has studied model similarity via internal representations, using SVCCA and CKA to show architectural correlations among models like BERT and GPT Wu et al. (2020). Conversely, Klabunde et al. (2023b) extended this analysis to 7B-scale models (LLaMA, Falcon, GPT-J), finding that representational similarity can vary even among similarly scaled or structured models. However, these methods are not applicable to proprietary models and do not capture behavioral traits like bias.
Alternative black-box methods compare decision boundaries Li et al. (2021) or prediction overlaps Guan et al. (2022), while some explore output-based similarity via accuracy or adversarial transferability Hwang et al. ; Jin et al. (2024). While efficient, scalar similarity metrics provide only a partial view, often leading to misinterpretation Klabunde et al. (2023a). This is especially problematic for generative models with large output spaces Klabunde et al. (2023b).
Our work reframes similarity through the lens of fairness—introducing bias similarity as a functional, behavior-based metric. We analyze 24 open- and closed-source LLMs, comparing how bias behaviors shift or persist across related models, highlighting family-level trends, tuning effects, and the limitations of superficial mitigation.
3 Bias Similarity Measurement Method
To analyze cross-model bias, we assess output distribution similarity across 24 LLMs. We define bias as disproportionate assumptions about certain groups, for instance, providing unbalanced answers favoring certain groups in ambiguous scenarios or consistently choosing stereotypical responses in forced-choice situations.
To measure bias similarities between LLMs, we input a prompt, consisting of a context, a question, and answer choices, to each model individually in a zero-shot fashion. We then collect the outputs and analyze their similarities using six metrics: accuracy, bias score, histogram, cosine distance, flip rates, and CKA. Each metric is computed based on models’ answer counts or output probabilities.
3.1 Models and Datasets
Models
We evaluated a diverse set of 24 LLMs from four prominent families:
  • LLaMA Family: Alpaca Taori et al. (2023), Vicuna Chiang et al. (2023), LLaMA 2 (7B, 7B-Chat) Touvron et al. (2023), LLaMA 3 (8B, 8B-Chat, 70B, 70B-Chat) Dubey et al. (2024).
  • Gemma Family: Gemma 1 (7B, 7B-It) Team et al. (2024a), Gemma 2 (9B, 9B-It, 27B, 27B-It) Team et al. (2024b), Gemma 3 (4B, 4B-It, 12B, 12B-It, 27B, 27B-It) Team et al. (2025).
  • GPT Family (API access): GPT-2 Radford et al. (2019), (as a baseline, though not directly comparable in scale/tuning to others), GPT-4o-mini 111platform.openai.com/docs/guides/text-generation (a state-of-the-art proprietary model).
  • Gemini Family (API access): Gemini-1.5-flash, Gemini-2.0-flash 222ai.google.dev/gemini-api/docs/models (representing another set of state-of-the-art proprietary models).
This selection includes both open-source and proprietary models, as well as base models and their instruction-tuned/chat-optimized variants, across a range of parameter sizes.
Datasets
We use three benchmark datasets for bias assessment: Bias Benchmark for QA (BBQ) Parrish et al. (2021), UnQover Li et al. (2020), and StereoSet Nadeem et al. (2020).
BBQ covers nine demographic dimensions, each with approximately 5K samples. Each sample includes a context, a question, and three multiple-choice answers (stereotype, anti-stereotype, and unknown), along with fairness-informed ground truth annotations. Crucially, it provides both ambiguous contexts (where “unknown” is the fairest answer) and disambiguated contexts (where sufficient information is provided to justify a specific correct answer). This setup allows us to evaluate whether models abstain appropriately (in ambiguous situations) and respond accurately when a definitive answer is required (in disambiguated setting).
@@ 3.1 Models and Datasets
UnQover is designed to uncover stereotypical bias through underspecified questions. It spans four demographic dimensions, each with 1M samples. Each sample consists of a context, a question, and two plausible answers, without ground-truth labels. Without an option to abstain, models are forced to reveal their biases through their preferences.
We focus on four dimensions—gender, race, religion, and nationality—across both datasets. Definitions and examples are provided in Appendix A.
We additionally evaluate bias through sentiment analysis in an open-ended generation task using a rephrased version of StereoSet. Details on this setup are described in Appendix G.
@@ 3.2 Similarity Assessment Metrics
We evaluate bias similarity using six complementary metrics that span performance, behavior, and internal representation: accuracy, bias score, output histograms, cosine distance, flip rates, and CKA.
Accuracy (BBQ Dismbiguated).
Each disambiguated BBQ question has a fairness-informed ground truth. We measure how often a model selects the correct, non-stereotypical response and avoids inappropriate abstention in contexts requiring a definitive answer.
Cosine Distance (UnQover and BBQ Ambiguous).
Cosine distance measures the angular difference between model output distributions across prompts, capturing alignment in relative preferences rather than absolute response frequencies Azarpanah and Farhadloo (2021). We represent each model’s response distribution as a vector and compute pairwise cosine similarity. Low cosine distance suggests proportional favoring of answers remains consistent, even if the absolute counts differ. Jensen–Shannon divergence is reported for comparison in Appendix E.
Bias Score (BBQ).
We adopt the bias score from Parrish et al. (2021) to quantify directional bias. Scores are defined separately for ambiguous and disambiguated contexts. [^3] The bias score for a disambiguated context question is defined as $s_{DIS}=2\left(\frac{n_{biased}}{n_{non\_unknown}}\right)-1$, where $n_{biased}$ and $n_{non\_unknown}$ refer to the number of biased and non-“unknown” answers, respectively. The score for an ambiguous question is defined as $s_{AMB}=(1-\text{acc})s_{DIS}$, where acc is the prediction accuracy on ambiguous questions. Values near 0 indicate neutrality, while $\pm$100 reflect bias toward stereotypes and anti-stereotypes, respectively.
Unknown (UNK) Flip Rates (BBQ Ambiguous)
For each base–tuned model pair, we measure the proportion of biased (stereotypical or anti-stereotypical) responses flipped to “Unknown” by the instruction-tuned model. Higher UNK flip rates indicate greater abstention in ambiguous cases, a desirable fairness behavior in this context.
Histogram (UnQover and BBQ).
@@
We visualize model outputs on UnQover and ambiguous BBQ prompts. Histograms reveal whether a model systematically favors certain responses, helping to identify underlying bias trends that scalar metrics may overlook.
Centered Kernel Alignment
CKA quantifies representational similarity between models by comparing their activation Gram matrices Kornblith et al. (2019). High CKA scores indicate that the representations learned by two models (or layers) are linearly transformable into one another, suggesting functional similarity. We compute CKA on penultimate-layer embeddings to assess alignment between base and instruction-tuned models.
Together, these metrics provide a multi-faceted view of how bias manifests across model outputs and internal representations.

### v4

# Bias Similarity Measurement: A Black-Box Audit of Fairness Across LLMs

@@ Abstract
Large Language Models (LLMs) reproduce social biases, yet prevailing evaluations score models in isolation, obscuring how biases persist across families and releases. We introduce Bias Similarity Measurement (BSM), which treats fairness as a relational property between models, unifying scalar, distributional, behavioral, and representational signals into a single similarity space. Evaluating 30 LLMs on 1M+ prompts, we find that instruction tuning primarily enforces abstention rather than altering internal representations; small models gain little accuracy and can become less fair under forced choice; and open-weight models can match or exceed proprietary systems. Family signatures diverge: Gemma favors refusal, LLaMA 3.1 approaches neutrality with fewer refusals, and converges toward abstention-heavy behavior overall. Counterintuitively, Gemma 3 Instruct matches GPT-4–level fairness at far lower cost, whereas Gemini’s heavy abstention suppresses utility. Beyond these findings, BSM offers an auditing workflow for procurement, regression testing, and lineage screening, and extends naturally to code and multilingual settings. Our results reframe fairness not as isolated scores but as comparative bias similarity, enabling systematic auditing of LLM ecosystems. Code is available at https://github.com/HyejunJeong/bias_llm.
@@ 1 Introduction
As AI systems increasingly influence societal decision-making in domains such as employment, finance, and law, ensuring model fairness has become a critical challenge to prevent adverse outcomes for protected groups (Ferrara, 2023). Large Language Models (LLMs) heighten this risk: they generate persuasive, human-like content at scale, and can reproduce or amplify social biases in sensitive contexts such as journalism, education, and healthcare (Sweeney, 2013). While many studies document biased behavior in individual models, we still lack a principled way to understand how such biases align, diverge, or persist across models and releases.
Bias in LLMs has been conceptualized in multiple ways: as systemic disparities across groups (Manvi et al., 2024), skewed performance across sociodemographic categories (Oketunji et al., 2023; Gupta et al., 2024), representational harms through stereotyping (Lin et al., 2025; Zhao et al., 2023), or unequal outcomes rooted in structural power imbalances (Gallegos et al., 2024). Yet defining bias remains nontrivial, since the line between bias and genuine demographic reflection is often blurred. For instance, if an LLM answers “younger people” to the question “Who tends to adapt to new technologies more easily: older or younger people?”, the response may be factually grounded in cognitive science (Vaportzis et al., 2017) but nonetheless reinforces stereotypes. This ambiguity motivates our study: rather than evaluating only scalar scores, we also analyze patterns of responses and abstentions, treating bias as a functional signature that can be compared across models.
Prior studies using fairness benchmarks, such as BBQ (Parrish et al., 2022), StereoSet (Nadeem et al., 2021), and UnQover (Li et al., 2020), assess models in isolation and provide scalar metrics, including bias scores or accuracy. While these reveal vulnerabilities, they provide no tools to analyze relationships between models. This omission matters: if fairness failures are structurally inherited, merely swapping one model for another may not resolve the problem. Conversely, if tuning strategies drive families toward convergent behaviors, then fairness gains may be superficial rather than structural. Without relational analysis, fairness audits risk overstating progress and underestimating systemic persistence.
We introduce Bias Similarity Measurement (BSM), a framework that treats bias as a relational property between models rather than an isolated attribute. Unlike prior work that analyzes models in isolation, BSM builds on functional similarity analyses (Klabunde et al., 2025; Li et al., 2021; Guan et al., 2022) but centers fairness as the dimension of comparison. Instead of asking “Is model $M$ biased?”, we ask, “Which models behave similarly with respect to bias, and why?” BSM integrates complementary similarity functions—scalar (accuracy, bias scores), distributional (histograms, cosine distance), behavioral (abstention flips), and representational (CKA)—into a unified space.
This reframing enables principled comparison across black-box systems and supports analyses not possible with prior metrics, such as detecting hidden lineage, quantifying family-level convergence, and tracking fairness drift across releases. It also grounds practical auditing tasks: procurement (balancing fairness and utility at abstention thresholds), regression testing (monitoring shifts across versions), and lineage screening (flagging suspiciously close bias profiles in proprietary systems).
Our evaluation covers 30 LLMs from four families (LLaMA, Gemma, GPT, Gemini), spanning 3B to 70B parameters, base and instruction-tuned variants, and both open- and closed-source systems. We analyze over 1M structured prompts from BBQ (Parrish et al., 2022) and UnQover (Li et al., 2020), plus open-ended generations from StereoSet (Nadeem et al., 2021). To our knowledge, this is the most comprehensive study of fairness similarity to date.
Contributions. Our contributions are threefold:
  • Conceptual/Methodological: We introduce BSM, a unified framework that reframes fairness as relational across models by integrating scalar, distributional, behavioral, and representational signals. This enables analyses not possible before, including lineage detection, family convergence, and fairness drift audits.
  • Empirical: We conduct the largest fairness similarity study to date—30 models from four families over 1M prompts—showing that fairness is dimension-specific and structurally uneven, defying capture by single scores.
  • Findings and Implications: Our findings include: instruction tuning enforces abstention rather than altering representations; on small models, tuning yields little gain and can reduce fairness; open models can match or exceed proprietary ones.
Motivating Example. Consider a start-up choosing a model for a customer-support assistant. Proprietary systems like GPT-4 or Gemini promise strong performance but at a high cost and with limited transparency. Open-weight options like Gemma 3-Instruct and LLaMA 3.1-Chat are more accessible and customizable, yet it is unclear which offers better fairness or utility. BSM provides a reusable evidence-based decision workflow, allowing practitioners to compare candidate models under fairness–utility constraints, rather than relying on reputation or size alone.
Figure 1: BSM Pipeline.
@@ 2 Related Works
We review two areas most relevant to our study: (i) how biases in LLMs have been evaluated, and (ii) how similarity across models has been assessed.
@@ 2.1 Bias Assessment in LLMs
Numerous studies have demonstrated that LLMs encode and reproduce social biases across various demographic dimensions. Early benchmarks such as StereoSet (Nadeem et al., 2021), CrowS-Pairs (Nangia et al., 2020), UnQover (Li et al., 2020), and BBQ (Parrish et al., 2022) introduced structured probes designed to expose stereotypical associations in templated or QA-style settings. More recent efforts broaden this space: CEB (Wang et al., 2025) and BEATS (Abhishek et al., 2025) expand coverage to multiple bias types and modalities, while Chaudhary et al. (2025) proposed formal certification of counterfactual bias. Other benchmarks, such as FairMT-Bench (Fan et al., 2025), move toward interactive multi-turn dialogue. Beyond datasets, LLMs themselves have been used as evaluators (Shi et al., 2024; Ye et al., 2025), though questions remain about consistency and induced bias (Stureborg et al., 2024). Architectural factors have also been studied (Yeh et al., 2023), as well as stereotype frequency (Bahrami et al., 2024) and retrieval exposure (Dai et al., 2024). Large-scale analysis by Kumar et al. (2024) evaluated implicit bias in 50+ models, finding that newer or larger models are not necessarily less biased and that provider-specific variation remains substantial. Despite this breadth, most prior work treats fairness as a property of individual models, reported as scalar metrics such as bias scores or accuracy. While these scores highlight vulnerabilities, they provide a siloed view of fairness behavior and do not capture how biases propagate across model families, scales, or tuning strategies. Our work instead bridges bias assessment and similarity analysis, reframing fairness as a relational property by comparing bias signatures across 30 open- and closed-source models.
@@ 2.2 LLM Similarity and Behavioral Alignment
In parallel, another line of work investigates similarity across models. At the representation level, SVCCA and CKA analyses reveal strong within-family correlations (Wu et al., 2020), although later studies note divergence across models of similar scale, such as LLaMA, Falcon, and GPT-J (Klabunde et al., 2025). Direct parameter comparisons, however, are often infeasible due to black-box APIs, architectural heterogeneity, and task mismatch (Li et al., 2021). To address this, black-box alternatives have been developed, comparing prediction overlaps (Guan et al., 2022), decision boundaries (Li et al., 2021), or adversarial transferability (Hwang et al., 2025; Jin et al., 2024). More recently, pipelines such as Polyrating (Dekoninck et al., 2025) introduce statistical rating schemes that account for evaluator biases (e.g., length or position effects) and align judgments across diverse tasks. While Polyrating incorporates fairness as one evaluation axis, its primary aim is comprehensive model scoring rather than dedicated analysis of fairness propagation. Thus, even when fairness is included, most similarity work does not place it at the center: they quantify alignment of representations or predictions, but not whether models replicate one another’s biases. We instead reframe similarity through fairness, introducing bias similarity as a functional, behavior-based metric that captures whether fairness patterns persist across families, regress across versions, or converge through alignment strategies such as abstention.
3 Bias Similarity Measurement
To answer the question, “How do LLMs exhibit biases across models?”, we introduce BSM, a framework that treats bias as a functional similarity relation between models rather than a fixed attribute of any single system. As illustrated in Figure 1, BSM systematically compares how multiple models behave under the same bias-sensitive prompts, by generating a bias similarity signature defined by four categories (categorical, distributional, behavioral, representational). The motivation is practical: with each new release claiming fairness improvements, what matters is not only the absolute bias level but also whether its bias profile inherits from, diverges from, or converges toward earlier versions and competing families.
3.1 Conceptual Framing
BSM interprets bias as a relational property emerging from comparing model outputs across demographic dimensions. We consider a set of models $\mathcal{M}=\{M_{1},\ldots,M_{n}\}$ and a set of bias dimensions $\mathcal{D}=\{d_{1},\ldots,d_{k}\}$ such as gender, race, nationality, and religion. Each dataset $\mathcal{X}$ consists of prompts $p\in\mathcal{X}$, where every prompt includes a context, a question, and a set of candidate answers. For a given model $M_{i}$, the predicted distribution on $p$ is denoted $M_{i}(p)$.
We define a bias similarity signature for each pair of models $(M_{i},M_{j})$ as a six-dimensional vector:
$\mathbf{S}(M_{i},M_{j}\mid\mathcal{X},\mathcal{D})=\big(S_{m_{1}},S_{m_{2}},\ldots,S_{m_{6}}\big),$
Each metric $S_{m_{\ell}}$ maps responses into a comparable form (categorical predictions, abstention markers, output distributions, or hidden representations) and computes similarity on distinct metrics (e.g., accuracy, bias score, cosine distance, histogram, flip rates, and CKA). Taken together, the signature provides a unified lens for comparing bias behaviors across models and dimensions.
3.2 Evaluation Pipeline
All models are evaluated on the same structured prompts spanning the bias dimensions in $\mathcal{D}$. Outputs are standardized: completions mapped to categorical labels, abstentions detected, distributions aggregated, and embeddings extracted where needed. Similarity functions $f_{m}$ are then applied pairwise to construct matrices summarizing bias similarity across the full model set. These matrices can be analyzed locally (within-family, e.g., base vs. tuned) or globally (e.g., open vs. proprietary), enabling comparisons of inheritance, divergence, and convergence across the ecosystem.
Metric Instantiations. Each metric captures a different facet of bias similarity. Accuracy on disambiguated questions evaluates whether two models converge on fairness-critical ground truth answers. Bias scores quantify directional skew in categorical predictions, revealing tendencies toward stereotypical or anti-stereotypical responses. Distributional comparisons, such as histograms and cosine distances, assess whether models allocate probability mass to answer categories in similar proportions. Abstention behavior is captured through unknown-flip rates, which measure whether biased answers are replaced by “Unknown.” Finally, CKA quantifies representational similarity, asking whether models encode prompts in linearly related feature spaces. Together, these instantiations span behavioral, functional, and representational levels of comparison.
Why a Unified Framework? Fairness evaluations often report a fragmented set of metrics, leaving it unclear how they relate to one another or whether they capture the same underlying mechanisms. BSM integrates behavioral, distributional, and representational measures into a single framework. This unification enables us to distinguish surface-level fairness behaviors from structural invariances, revealing, for example, that instruction tuning may leave representational bias intact while enforcing behavioral convergence through abstention. This integrative perspective is essential for developing robust fairness audits in an ecosystem where models are diverse, fast-evolving, and often only accessible as black-box APIs.
Scope and inference. We distinguish controlled, within-family comparisons (same base weights; tuning is the primary difference), which support interpretive claims about instruction-tuning effects, from cross-vendor comparisons (architecture/data/pipelines differ), which we report as observational ecosystem mapping only. We avoid causal language for cross-vendor results and report uncertainty for within-family deltas.
4 Evaluation Setup
4.1 Models
We evaluated a diverse set of 30 LLMs from four families: LLaMA: Vicuna (Chiang et al., 2023), LLaMA 2 (7B) (Touvron et al., 2023), LLaMA 3/3.1 (8B, 70B) (Dubey et al., 2024), and LLaMA 3.2 (4B), each with -Chat variants (Meta AI, 2024). Gemma: Gemma 1 (7B), Gemma 2 (9B, 27B), and Gemma 3 (4B, 12B, 27B), each with -It variants  (Team et al., 2024a; b; 2025). GPT: GPT-2 (Radford et al., 2019), (as a baseline), GPT-4o-mini (OpenAI, 2024), and GPT-5-mini (OpenAI, 2025) [^1] We exclude GPT-5-mini from UnQover due to the prohibitive cost of running it across the full sample set.. Gemini: Gemini-1.5-flash and 2.0-flash (Google AI Developers, 2025).
The “-Chat” or “-It” suffixes denote instruction-tuned variants, optimized for conversational use and typically exhibiting fewer safety violations (Touvron et al., 2023). Our selection spans open-source and proprietary models, base and instruction-tuned variants, and multiple parameter scales, enabling comparisons both across and within families.
4.2 Datasets
We use three complementary benchmarks: BBQ (Parrish et al., 2022), UnQover (Li et al., 2020), and StereoSet (Nadeem et al., 2021) to cover fairness-labeled, forced-choice, and open-ended settings.
BBQ spans nine demographic dimensions with $\sim$5K samples each. Each prompt includes a context, question, and three answers (stereotype, anti-stereotype, unknown), with fairness-informed ground truth. Ambiguous contexts make “unknown” the fairest option, while disambiguated contexts require a definitive answer, enabling evaluation of abstention vs. accuracy.
@@ 4.2 Datasets
UnQover probes bias through underspecified questions across four dimensions ($\sim$1M samples). Each consists of a context, question, and two plausible answers, without ground truth or abstention, forcing models to reveal directional bias.
We align our analysis on the four dimensions common to both (gender, race, religion, nationality), with definitions in Table 4. We also extend to open-ended generation via a rephrased StereoSet, detailed in Appendix H.
@@ 4.3 Similarity Assessment Metrics
To capture the multifaceted nature of bias similarity, we evaluate models with six complementary metrics spanning accuracy, behavioral tendencies, output distributions, and internal representations.
Accuracy (BBQ Dismbiguated). Each disambiguated BBQ question has a ground truth answer indicating fairness. We use accuracy to measure functional similarity between LLMs, reflecting both fairness and contextual understanding. In disambiguated contexts, where the correct answer is clear given sufficiently informative context, accuracy reveals whether bias overrides correct choices.
Unknown (UNK) Flip Rates (BBQ Ambiguous). For each base–tuned model pair, we introduce UNK Flip as a pairwise measure of abstention shifts under instruction tuning. For a base model $M_{b}$ and tuned model $M_{t}$, it is defined as
$\text{UNK Flip}(M_{b}\!\to\!M_{t})=\frac{n_{\text{biased}\to\text{UNK}}}{n_{\text{biased}}},$
where $n_{\text{biased}}$ is the number of biased responses (stereotypical or anti-stereotypical) from $M_{b}$, and $n_{\text{biased}\to\text{UNK}}$ is the subset flipped to “Unknown” by $M_{t}$. High values indicate that tuning promotes abstention in underspecified contexts, mitigating bias reinforcement, while low values suggest limited fairness gains.
Bias Score (BBQ). We adopt the bias score from (Parrish et al., 2022) to quantify directional bias, defined separately depending on question contexts. The scores are defined as follows:
$s_{\text{DIS}}$ $=2\left(\frac{n_{\text{biased}}}{n_{\text{non\_unknown}}}\right)-1,$ $\quad s_{\text{AMB}}$ $=(1-\text{acc})\,s_{\text{DIS}}.$
Here $n_{\text{biased}}$ and $n_{\text{non\_unknown}}$ are the counts of biased and non-“unknown” responses, and acc is the accuracy on ambiguous questions. We report scores multiplied by 100 for readability, so values range from 100-100 (anti-stereotypical) to +100+100 (stereotypical), with near 0 indicating neutrality.
@@
Histogram (UnQover and BBQ Ambiguous). Although accuracy and bias scores quantify performance, allowing a convenient comparison across models, they do not reveal distributional patterns. We therefore visualize model outputs on UnQover and ambiguous BBQ prompts. Histograms reveal whether a model systematically favors certain responses, identifying underlying bias trends that scalar metrics may overlook.
Cosine Distance (UnQover and BBQ Ambiguous). We use cosine distance to compare model output distributions across prompts, following prior work on count-based similarity measures (Azarpanah & Farhadloo, 2021; Singhal et al., 2017; Kocher & Savoy, 2017). Unlike scalar accuracy, cosine distance captures alignment in relative preferences rather than absolute frequencies. We compute distances directly on raw count vectors (without normalization), so low values indicate stable proportional preferences even if absolute counts differ. For completeness, we report Jensen–Shannon divergence results in Appendix G.
Centered Kernel Alignment (CKA). CKA measures representational similarity by comparing activation patterns (i.e., Gram matrices) across models (Kornblith et al., 2019). Unlike output-based metrics, it probes internal feature spaces: high scores indicate that models encode inputs in linearly related ways, suggesting structural similarity even if outputs differ. In our setting, CKA examines how instruction tuning affects internal representations and whether representational similarity correlates with changes in output behavior, thereby clarifying whether tuning alters reasoning pathways or primarily impacts surface responses.
Together, these metrics capture both the magnitude and structure of bias, offering a balanced view of performance, behavior, and representations, and enabling a comprehensive assessment of how instruction tuning, version increments, and institutional differences shape outputs and internal mechanisms across families and scales.
Table 1: Average bias scores. “–”: anti-stereotypical, “+”: stereotypical, and values near 0 = neutrality. Shown for ambiguous (s_AMB) and disambiguated (s_DIS) contexts.
Base Model Avg. s_AMB Avg. s_DIS
Base Tuned Base Tuned
LLaMA 2 7B 5.45 4.30 7.50 6.65
LLaMA 3 8B -4.78 -0.66 -8.72 -2.10
LLaMA 3 70B -1.42 0.55 -4.51 2.50
LLaMA 3.1 8B 18.59 1.38 31.37 4.78
LLaMA 3.1 70B 0.42 -0.15 0.81 -1.44
LLaMA 3.2 3B 11.95 15.71 17.67 30.97
Gemma 7B 1.81 2.05 0.69 0.95
Gemma 2 9B 0.08 0.18 6.83 -2.02
Gemma 2 27B 6.95 0.51 14.31 -1.45
Gemma 3 4B -3.89 5.83 2.69 8.62
Gemma 3 12B 4.36 0.15 6.12 -0.17
Gemma 3 27B -1.25 0.07 -0.26 -1.50
Gemini 1.5 2.37 3.07
Gemini 2.0 -4.17 -5.54
GPT-2 72.43 96.19
GPT-4o Mini 0.47 2.66
GPT-5 Mini 0.21 1.10